# SReferences

### v2

@@ 1 Introduction
Market efficiency is a fundamental concept in econometrics, which indicates how fast the market can absorb new information and how swiftly market prices change to reflect the fair value of an asset. The Efficient Market Hypothesis, first introduced by Fama, (1970), states that, in an ideal situation, the market price of an asset reflects all available information in the market and must be equal to the asset’s fair value. In other words, when the market is fully efficient, the price returns are entirely unpredictable. The debate on the use of the Efficient Market Hypothesis for statistical analysis has been ongoing for decades, and the Efficient Market Hypothesis has been challenged by many empirical studies; see, e.g., Lo, (2004, 2005) and Malkiel, (2003). Various statistical methods have been proposed to test the Efficient Market Hypothesis (Lo and MacKinlay,, 1988; Choi,, 1999; Escanciano and Lobato,, 2009; Charles et al.,, 2011; Wang,, 2024). However, these methods are often based on the assumption that log-returns are normally distributed, while, in reality, the returns typically have much heavier tails and non-Gaussian dependencies, especially in a high-frequency setting (Cont,, 2001; De Domenico et al.,, 2023). Moreover, increasing evidence shows that the market is not always efficient, especially when it experiences extreme events during financial turmoil (Choi,, 2021); this kind of inefficiency in tails is often ignored by traditional Efficient Market Hypothesis tests.
@@
We apply our method to study the market-wide tail-efficiency of China’s futures market, using high-frequency data containing second-level market information that we mined for use in statistical analyses. In our analysis, we find that most of the time-leading tail dependencies are symmetric, i.e., $\lambda(X_{i},Y_{j})=0,$ but some inefficiencies persist. To capitalize on this finding, we construct an artificial dynamic portfolio consisting of assets that do not satisfy the conditions of the ETH and demonstrate that it can generate profits using out-of-sample data.
@@
The rest of the paper is organized as follows: Section 2 reviews the literature on regular variation and the EDM. Section 3 introduces balanced regular variation, our directional tail dependence measure, the ETH, and our hypothesis test. Section 4 provides a simulation study that gives empirical evidence for the power of our permutation test. Section 5 presents our application to asset returns from China’s futures market; the dataset and its construction are described in the Supplementary Material, Section B. Relevant proofs are included in the Supplementary Material, Section A. Supporting code and the dataset used in our analysis are available at the Github repository, https://github.com/Stathastic/tailEfficientTest. The full dataset can be obtained upon request.
@@ 2 Multivariate extreme-value background
where $\oplus$ and $\circ$ are transformed-linear operators (see Cooley and Thibaud,, 2019, for details) that satisfy $X_{1}\oplus X_{2}=t\{t^{-1}(X_{1})+t^{-1}(X_{2})\}$ and $a\circ X=t\{at^{-1}(X)\}$, for constant $a\in\mathbb{R}$ and ${t(\cdot)=\log\{1+\exp(\cdot)\}}$. The space $\mathcal{V}^{q}_{+}$ is a subspace of ${\rm RV}^{1}_{+}(2)$, and contains elements that can be spanned (using the transformed-linear operations) by a vector of independent regularly varying ${\rm RV}^{1}_{+}(2)$ random variables with non-negative coefficients. This space has two important properties, which make it conducive to studying the tails of random variables (see, e.g., the transformed linear model of Mhatre, (2022) and the partial tail correlation coefficients of Lee and Cooley, (2022) and Gong et al., (2024)): first, the corresponding angular measure $H_{\mathbf{X}}(\cdot)$ for any $\bm{X}\in{\rm RV}^{p}_{+}(2)$ can be approximated using a sequence of transformed-linear combinations of independent regularly varying random variables (Proposition 4; Cooley and Thibaud,, 2019) with increasing $q$. This implies that $\mathcal{V}^{q}_{+}$ is dense in ${\rm RV}^{p}_{+}(2)$ as $q\to\infty$. Second, the inner product in $\mathcal{V}^{q}_{+}$, i.e., $\langle X,Y\rangle:=\bm{a}_{x}^{T}\bm{a}_{y}$ where $\bm{a}_{x}$ and $\bm{a}_{y}$ are the coefficient vectors of $X$ and $Y$, respectively, in the transformed-linear space $\mathcal{V}^{q}_{+}$, can be related to the Extremal Dependence Measure (EDM). This measure was first introduced by Resnick, (2004) and later studied by Larsson and Resnick, (2012) and Cooley and Thibaud, (2019).
@@ 3.1 Balanced regular variation
Let upper tail $X_{i}^{+}:=\max\{X_{i},0\}$ and lower tail $X_{i}^{-}:=-\min\{X_{i},0\}$ for $i=1,\dots,p$. Then ${\bm{X}^{+}:=(X_{1}^{+},\dots,X_{p}^{+})^{\top}}$ and ${\bm{X}^{-}:=(X_{1}^{-},\dots,X_{p}^{-})^{\top}}$ contain information on the upper and lower tails, respectively, of ${\bm{X}\in{\rm BRV}^{p}(2)}$. We concatenate $\bm{X}^{+}$ and $\bm{X}^{-}$ into the $2p$-dimensional vector $\bar{\bm{X}}:=\left((\bm{X}^{+})^{\top},(\bm{X}^{-})^{\top}\right)^{\top}$. Properties of $\bar{\bm{X}}$ are described in Proposition 3.2 (see the Supplementary Material, Section A.2 for the proof).
@@ 3.2 Marginal transformation to balanced regular variation
In practice, the upper and lower tails of a random variable may exhibit different tail heaviness, and the margins of our asset return data, denoted here by $\mathbf{R}\in\mathbb{R}^{p},$ may not be balanced, i.e., they may have different scales or tail-heaviness. However, the methodology in this paper remains applicable after standardizing the data margins to ensure balanced regular variation is satisfied. The pre-standardization of data to common margins is common practice for modeling of multivariate extremes (see discussion by, e.g., Naveau and Segers,, 2024). We describe here two marginal transformations of $\mathbf{R}$ to a balanced regularly varying random vector ${\rm BRV}^{p}(2)$: the first assumes that the upper and lower tails of all components of $\mathbf{R}$ are marginally regularly varying (we find this to be the case for the data in our application; see Section 5 and Section D.2 of the Supplementary Material), whilst the second is applicable for general data margins.
@@ 3.3 Visualizing directional tail asymmetry
The two boundary cases of the tail efficient line, where ${\sigma(X^{+},Y^{+})=\sigma(X^{+},Y^{-})=\frac{\sqrt{2}}{4}}$ and ${\sigma(X^{+},Y^{+})=\sigma(X^{+},Y^{-})=0}$, are achieved when $X^{+}$ is at the most strongly dependent tail efficient point $(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0)$ and mutual asymptotic independence point $(0,0,1)$, respectively. The most strongly dependent tail efficient point indicates the upper limit of achievable extremal dependence between $X^{+}$ and $Y^{+}$ (or $X^{+}$ and $Y^{-}$), under directional symmetry. The mutual asymptotic independence point reflects another special case where $X^{+}$, $Y^{+}$, and $Y^{-}$ are all mutually asymptotically independent. The extremal dependence for $(X^{+},Y^{+})^{\top}$ and $(X^{+},Y^{-})^{\top}$ are entangled, as stated in Proposition 3.3 (see the Supplementary Material, Section A.3 for the proof).
@@ 3.4 Directional tail dependence measure
The DTD measure can be used to describe the dependence from target $Y$ on the explanatory variable $X$, particularly when focusing on extremal situations. The case $\lambda(X,Y)=0$ implies that the dependence structure for $(X^{+},Y^{+},Y^{-})^{\top}$ is directionally symmetric. We design this measure to satisfy certain desirable properties; see, e.g., Embrechts et al., (2002). Specifically, the DTD measure satisfies the following two properties as stated in Proposition 3.4 (see the Supplementary Material, Section A.4 for the proof).
@@
We can view the first form of $\lambda$ in Equation (7) as an extension of the EDM for the vector $(X^{+},Y)^{\top}$, which is regularly varying on the right half space of $\mathbb{R}^{2}$. The second form of $\lambda$ is the difference between the quadrant-specific measures $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$. These two forms are equivalent, but permit two different estimators; for the proof of Proposition 3.5, see the Supplementary Material, Section A.5.
@@
The estimator $\widehat{\lambda}^{1}_{n}(X,Y)$ is asymptotically normal under similar assumptions as the EDM; see Larsson and Resnick, (2012), Theorem 1, and Proposition 3.6 below (proved in the Supplementary Material, Section A.6).
@@
Asymptotic properties for $\widehat{\lambda}^{2}_{n}$ are more difficult to derive and, instead, we construct a permutation test based on this statistic to check the Efficient Tail Hypothesis; see Section 3.5.
@@ 3.5 Efficient Tail Hypothesis (ETH)
Testing of the ETH for vectors $\bm{X}$ and $\bm{Y}$ requires multiple tests, which increases the risk of obtaining false positive results, i.e., incorrectly rejecting a null hypothesis. Many methods have been proposed to correct the multiple testing problem (see, e.g., Bonferroni,, 1936; Holm,, 1979; Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001). In Section 5, we use the Benjamini–Hochberg correction to test for tail efficiency of China’s futures market.
@@ 4 Simulation study
Here we investigate the power of the tail efficiency tests for bivariate random vectors. Section 4.1 introduces the simulation design and Section 4.2 presents the results of the study.
@@ 4.1 Simulation design
Figure 2: Scatter plots for simulated $(U,V)^{\top}$ under the null hypothesis with $\phi=0.5$ (Left two columns) and alternative hypothesis with $\phi=0.7$ (Right two columns). The underlying copula differs with the columns: the first and third columns are scatter points from the Gumbel copula and the second and fourth columns are scatter points from the Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1,1.2,1.5,2\}$; for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.
@@ 4.2 Results
For each setting, we test the null $\mathcal{H}_{0}:\lambda(X,Y)=0$ for $K=1,000$ repetitions. For each test, we simulate independent samples $\{(x_{i},y_{i})^{\top}\}_{i=1}^{n}$ with sample size $n=10,000$ and calculate the test statistic with threshold $r_{0}$ taken to be the empirical $0.99$-quantile of $\{r_{i}=||(x_{i},y_{i})^{\top}||_{2}\}_{i=1}^{n}$ (and similarly for $r_{0}^{+}$ and $r_{0}^{-}$). We compare the permutation test (with $P=1,000$) for both estimators $\widehat{\lambda}^{1}_{n}$ and $\widehat{\lambda}^{2}_{n}$ in Definition 3.5, with the $t$-test based on the asymptotic normality of $\widehat{\lambda}^{1}_{n}$. In the end, for each of the three test types, we have $K$ independent pairs of test statistics $|\widehat{\lambda}_{n}^{(k)}|$ and their corresponding p-value, $p^{(k)}$, for $k=1,\dots,K$. Figure 3 visualises the relationship between the test statistics and their p-values.
@@
Figure 3: Plot of p-values ($y$-axis) against empirical rank of test statistic $|\widehat{\lambda}_{n}^{(k)}(X,Y)|/K$ ($x$-axis); case $\lambda(X,Y)=0$ (Left two columns) and case $\lambda(X,Y)\neq 0$ (Right two columns); Red and blue points correspond to permutation tests using $\widehat{\lambda}_{n}^{1}$ and $\widehat{\lambda}^{2}_{n}$, respectively; purple points correspond to $t$-tests using $\widehat{\lambda}_{n}^{1}$. The underlying copula differs with the columns: the first and third columns are results for the Gumbel copula and the second and fourth columns are results for the Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1,1.2,1.5,2\}$; for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.
@@
Under the null hypothesis, $p^{(k)}$ should be roughly equal to $1-{\rm rank}(|\widehat{\lambda}^{(k)}_{n})|/K$. In Figure 3, left column-block, we see that all points are clustered around the second diagonal under symmetry, which means that the rejection percentages for each significance level are approximately equal to the corresponding nominal value. From the right column-block of Figure 3, which displays the results when data are simulated under the alternative hypothesis (with $\phi=0.7$), all points are below the diagonal line. Points closer to the $x$-axis indicate tests with larger power of rejection. As expected, the statistical power for all three testing methods become stronger with increasing tail asymmetry (descending rows).
@@
The permutation test using the second estimator $\widehat{\lambda}^{2}_{n}$ appears to have the strongest power among the three testing methods. This may be attributed to its data efficiency, as $\widehat{\lambda}^{2}_{n}$ uses more unique data than $\widehat{\lambda}^{1}_{n}$. Recall from Definition 3.5 that $\widehat{\lambda}^{2}_{n}$ is an estimator for ${\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}}$, which involves estimating the EDM for $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$, separately. For each permuted samples, the set of indices for which the $l_{2}$-norm exceeds the threshold $r^{+}_{0}$ and $r_{0}^{-}$ will vary; for estimator $\widehat{\lambda}^{1}_{n}$, this set remains consistent across all $P$ permutations. The advantage of using $\widehat{\lambda}^{2}_{n}$ comes with a higher computational cost as estimation requires re-selecting the exceedances for each permutation. Another advantage of using $\widehat{\lambda}^{2}_{n}$ is that it has the flexibility of choosing two thresholds, $r_{0}^{+}$ and $r_{0}^{-}$, instead of a single threshold $r_{0}$.
We also investigate the effect of the sample size $n$ and lower exceedance thresholds (see the Supplementary Material, Section C); the results are similar, and so are not discussed further here. In our application in Section 5, we use the permutation-based test with estimator $\widehat{\lambda}^{2}_{n}$.

### v3

@@ 1 Introduction
Market efficiency is a fundamental concept in finance, which indicates how fast the market can absorb new information and how swiftly market prices change to reflect the fair value of an asset. The Efficient Market Hypothesis, first introduced by Fama, (1970), states that, in an ideal situation, the market price of an asset reflects all available information in the market and must be equal to the asset’s fair value. In other words, when the market is fully efficient, the price returns are entirely unpredictable. The debate on the use of the Efficient Market Hypothesis for statistical analysis has been ongoing for decades, and the Efficient Market Hypothesis has been challenged by many empirical studies; see, e.g., Lo, (2004, 2005) and Malkiel, (2003). Various statistical methods have been proposed to test the Efficient Market Hypothesis (Lo and MacKinlay,, 1988; Choi,, 1999; Escanciano and Lobato,, 2009; Charles et al.,, 2011; Wang,, 2024). However, these methods are often based on the assumption that log-returns are normally distributed, while, in reality, the returns typically have much heavier tails and non-Gaussian dependencies, especially in a high-frequency setting (Cont,, 2001; De Domenico et al.,, 2023). Moreover, increasing evidence shows that the market is not always efficient, especially when it experiences extreme events during financial turmoil (Choi,, 2021); this kind of inefficiency in tails is often ignored by traditional Efficient Market Hypothesis tests.
@@
We apply our method to study the market-wide tail-efficiency of China’s futures market, using high-frequency data containing second-level market information that we mined for use in statistical analyses. We find that most of the time-leading tail dependencies are symmetric, i.e., $\lambda(X_{i},Y_{j})=0,$ but some inefficiencies persist. To capitalize on this finding, we construct an artificial dynamic portfolio consisting of assets that do not satisfy the conditions of the ETH and show that it can generate profits using out-of-sample data.
@@
The rest of the paper is organized as follows: Section 2 reviews the literature on regular variation and the EDM. Section 3 introduces balanced regular variation, our directional tail dependence measure, the ETH, and our hypothesis test. Section 4 provides a simulation study that gives empirical evidence for the power of our permutation test. Section 5 presents our application to asset returns from China’s futures market; the dataset and its construction are described in the Appendix B. Relevant proofs are included in the Appendix A. Supporting code and the dataset used in our analysis are available at the Github repository, https://github.com/junshujiang/tailEfficientTest. The full dataset can be obtained upon request.
@@ 2 Multivariate extreme-value background
where $\oplus$ and $\circ$ are transformed-linear operators (see Cooley and Thibaud,, 2019, for details) that satisfy $X_{1}\oplus X_{2}=t\{t^{-1}(X_{1})+t^{-1}(X_{2})\}$ and $a\circ X=t\{at^{-1}(X)\}$, for constant $a\in\mathbb{R}$ and ${t(\cdot)=\log\{1+\exp(\cdot)\}}$. The space $\mathcal{V}^{q}_{+}$ is a subspace of ${\rm RV}^{1}_{+}(2)$, and contains elements that can be spanned (using the transformed-linear operations) by a vector of independent regularly varying ${\rm RV}^{1}_{+}(2)$ random variables with non-negative coefficients. This space has two important properties, which make it conducive to studying the tails of random variables (see, e.g., the transformed linear model of Mhatre, (2022) and the partial tail correlation coefficients of Lee and Cooley, (2022) and Gong et al., (2024)): first, the corresponding angular measure $H_{\mathbf{X}}(\cdot)$ for any $\bm{X}\in{\rm RV}^{p}_{+}(2)$ can be approximated using a sequence of transformed-linear combinations of independent regularly varying random variables (Proposition 4; Cooley and Thibaud,, 2019) with increasing $q$. This implies that $\mathcal{V}^{q}_{+}$ is dense in ${\rm RV}^{1}_{+}(2)$ as $q\to\infty$. Second, the inner product in $\mathcal{V}^{q}_{+}$, i.e., $\langle X,Y\rangle:=\bm{a}_{x}^{T}\bm{a}_{y}$ where $\bm{a}_{x}$ and $\bm{a}_{y}$ are the coefficient vectors of $X$ and $Y$, respectively, in the transformed-linear space $\mathcal{V}^{q}_{+}$, can be related to the Extremal Dependence Measure (EDM). This measure was first introduced by Resnick, (2004) and later studied by Larsson and Resnick, (2012) and Cooley and Thibaud, (2019).
@@ 3.1 Balanced regular variation
Let upper tail $X_{i}^{+}:=\max\{X_{i},0\}$ and lower tail $X_{i}^{-}:=-\min\{X_{i},0\}$ for $i=1,\dots,p$. Then ${\bm{X}^{+}:=(X_{1}^{+},\dots,X_{p}^{+})^{\top}}$ and ${\bm{X}^{-}:=(X_{1}^{-},\dots,X_{p}^{-})^{\top}}$ contain information on the upper and lower tails, respectively, of ${\bm{X}\in{\rm BRV}^{p}(2)}$. We concatenate $\bm{X}^{+}$ and $\bm{X}^{-}$ into the $2p$-dimensional vector $\bar{\bm{X}}:=\left((\bm{X}^{+})^{\top},(\bm{X}^{-})^{\top}\right)^{\top}$. Properties of $\bar{\bm{X}}$ are described in Proposition 3.2 (see the Appendix A.2 for the proof).
@@ 3.2 Marginal transformation to balanced regular variation
In practice, the upper and lower tails of a random variable may exhibit different tail heaviness, and the margins of our asset return data, denoted here by $\mathbf{R}\in\mathbb{R}^{p},$ may not be balanced, i.e., they may have different scales or tail-heaviness. However, the methodology in this paper remains applicable after standardizing the data margins to ensure balanced regular variation is satisfied. The pre-standardization of data to common margins is common practice for modeling of multivariate extremes (see discussion by, e.g., Naveau and Segers,, 2024). We describe here two marginal transformations of $\mathbf{R}$ to a balanced regularly varying random vector ${\rm BRV}^{p}(2)$: the first assumes that the upper and lower tails of all components of $\mathbf{R}$ are marginally regularly varying (we find this to be the case for the data in our application; see Section 5 and the Appendix D.2), whilst the second is applicable for general data margins.
@@ 3.3 Visualizing directional tail asymmetry
The two boundary cases of the tail efficient line, where ${\sigma(X^{+},Y^{+})=\sigma(X^{+},Y^{-})=\frac{\sqrt{2}}{4}}$ and ${\sigma(X^{+},Y^{+})=\sigma(X^{+},Y^{-})=0}$, are achieved when $X^{+}$ is at the most strongly dependent tail efficient point $(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0)$ and mutual asymptotic independence point $(0,0,1)$, respectively. The most strongly dependent tail efficient point indicates the upper limit of achievable extremal dependence between $X^{+}$ and $Y^{+}$ (or $X^{+}$ and $Y^{-}$), under directional symmetry. The mutual asymptotic independence point reflects another special case where $X^{+}$, $Y^{+}$, and $Y^{-}$ are all mutually asymptotically independent. The extremal dependence for $(X^{+},Y^{+})^{\top}$ and $(X^{+},Y^{-})^{\top}$ are entangled, as stated in Proposition 3.3 (see the Appendix A.3 for the proof).
@@ 3.4 Directional tail dependence measure
The DTD measure can be used to describe the dependence from target $Y$ on the explanatory variable $X$, particularly when focusing on extremal situations. The case $\lambda(X,Y)=0$ implies that the dependence structure for $(X^{+},Y^{+},Y^{-})^{\top}$ is directionally symmetric. We design this measure to satisfy certain desirable properties; see, e.g., Embrechts et al., (2002). Specifically, the DTD measure satisfies the following two properties as stated in Proposition 3.4 (see the Appendix A.4 for the proof).
@@
We can view the first form of $\lambda$ in Equation (7) as an extension of the EDM for the vector $(X^{+},Y)^{\top}$, which is regularly varying on the right half space of $\mathbb{R}^{2}$. The second form of $\lambda$ is the difference between the quadrant-specific measures $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$. These two forms are equivalent, but permit two different estimators; for the proof of Proposition 3.5, see the Appendix A.5.
@@
The estimator $\widehat{\lambda}^{1}_{n}(X,Y)$ is asymptotically normal under similar assumptions as the EDM; see Larsson and Resnick, (2012), Theorem 1, and Proposition 3.6 below (proved in the Appendix A.6).
@@
[…]
@@ 3.5 Efficient Tail Hypothesis (ETH)
In the application to financial markets (Section 5), we let the target random variables, contained within $\bm{Y},$ denote financial asset returns, and the explanatory variables, in $\bm{X}$, represent available prior information available. The null hypothesis of the ETH states that the explanatory variables exhibit symmetric tail dependence with respect to both the upper and lower tails of asset returns. When the ETH is rejected, it suggests that the explanatory variables can help to predict the direction of large price movements in the target assets, indicating market inefficiency. However, when the ETH cannot be rejected, one cannot conclude that the market is efficient, not only because statistical tests can never be used to formally confirm a hypothesis when it is true—they can only provide evidence against it when it is false—but more importantly because the statistic $\lambda$ focuses only on extreme events.
@@
Testing of the ETH for vectors $\bm{X}$ and $\bm{Y}$ requires multiple tests, which increases the risk of obtaining false positive results, i.e., incorrectly rejecting a null hypothesis. Many methods have been proposed to correct the multiple testing problem (see, e.g., Bonferroni,, 1936; Holm,, 1979; Benjamini and Hochberg,, 1995). In Section 5, we use the Benjamini–Hochberg correction to test for tail efficiency of China’s futures market.
@@ 4 Simulation study
Here we investigate the power of the tail efficiency tests for bivariate random vectors. Section 4.1 introduces the simulation design and Section 4.2 presents the results.
@@ 4.1 Simulation design
Figure 2: Scatter plots for simulated $(U,V)^{\top}$ under the null hypothesis with $\phi=0.5$ (Left two columns) and alternative hypothesis with $\phi=0.7$ (Right two columns). The underlying copula differs with the columns: the first and third columns are scatter points from the Gumbel copula and the second and fourth columns are scatter points from the Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1.1,1.5,2.5,3.0\}$; for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.
@@ 4.2 Results
For each setting, we test the null $\mathcal{H}_{0}:\lambda(X,Y)=0$ for $K=1,000$ repetitions. For each test, we simulate independent samples $\{(x_{i},y_{i})^{\top}\}_{i=1}^{n}$ with sample size $n=10,000$ and calculate the test statistic with threshold $r_{0}$ taken to be the empirical $q=0.99$-quantile of $\{r_{i}=||(x_{i},y_{i})^{\top}||_{2}\}_{i=1}^{n}$ (and similarly for $r_{0}^{+}$ and $r_{0}^{-}$). We compare the permutation test (with $P=1,000$) for both estimators $\widehat{\lambda}^{1}_{n}$ and $\widehat{\lambda}^{2}_{n}$ in Definition 3.5, with the $t$-test based on the asymptotic normality of $\widehat{\lambda}^{1}_{n}$. In the end, for each of the three test types, we have $K$ independent pairs of test statistics $|\widehat{\lambda}_{n}^{(k)}|$ and their corresponding p-value, $p^{(k)}$, for $k=1,\dots,K$. Figure 3 visualises the relationship between the test statistics and their p-values. We also investigate the effect of the sample size $n$ and lower exceedance thresholds (see the Appendix C); the results are similar, and so are not discussed further here.
@@
Figure 3: Plot of p-values ($y$-axis) against empirical rank of test statistic $|\widehat{\lambda}_{n}^{(k)}(X,Y)|/K$ ($x$-axis); case $\lambda(X,Y)=0$ (Left two columns) and case $\lambda(X,Y)\neq 0$ (Right two columns); Circular (green) and triangular (yellow) points correspond to permutation tests using $\widehat{\lambda}_{n}^{1}$ and $\widehat{\lambda}^{2}_{n}$, respectively; Square (blue) points correspond to $t$-tests using $\widehat{\lambda}_{n}^{1}$. The underlying copula differs with the columns: the first and third columns are results for the Gumbel copula and the second and fourth columns are results for the Student’s $t$-copula. Rows correspond to different parameters for the copulas: for the Gumbel copula, $\theta\in\{1.1,1.5,2.5,3.0\}$; for the Student’s $t$-copula, $\rho\in\{0.1,0.4,0.6,0.8\}$ with degrees of freedom $\nu$ fixed to be $4$.
@@
Under the null hypothesis, $p^{(k)}$ should be roughly equal to $1-{\rm rank}(|\widehat{\lambda}^{(k)}_{n})|/K$. In Figure 3, left column-block, we see that all points are clustered around the second diagonal under symmetry, which means that the rejection percentages for each significance level are approximately equal to the corresponding nominal value (see the Appendix C for further analysis on the empirical sizes of the three tests). From the right column-block of Figure 3, which displays the results when data are simulated under the alternative hypothesis (with $\phi=0.7$), all points are below the diagonal line. Points closer to the $x$-axis indicate tests with larger power of rejection.
@@
The power of the tests, estimated as the percentage of rejections of the null hypothesis, when data are generated under the alternative hypothesis is investigated in Table 1 and, as expected, the statistical power for all three testing methods becomes higher with increasing tail asymmetry (across descending rows in Figure 3). The permutation test using the second estimator $\widehat{\lambda}^{2}_{n}$ have the strongest power among the three testing methods. This may be attributed to its data efficiency, as $\widehat{\lambda}^{2}_{n}$ uses more unique data than $\widehat{\lambda}^{1}_{n}$. Recall from Definition 3.5 that $\widehat{\lambda}^{2}_{n}$ is an estimator for ${\lambda(X,Y)=2\{\sigma(X^{+},Y^{+})-\sigma(X^{+},Y^{-})\}}$, which involves estimating the EDM for $\sigma(X^{+},Y^{+})$ and $\sigma(X^{+},Y^{-})$, separately. For each permuted samples, the set of indices for which the $l_{2}$-norm exceeds the threshold $r^{+}_{0}$ and $r_{0}^{-}$ will vary; for estimator $\widehat{\lambda}^{1}_{n}$, this set remains consistent across all $P$ permutations. The advantage of using $\widehat{\lambda}^{2}_{n}$ comes with a higher computational cost as estimation requires re-selecting the exceedances for each permutation. Another advantage of using $\widehat{\lambda}^{2}_{n}$ is that it has the flexibility of choosing two thresholds, $r_{0}^{+}$ and $r_{0}^{-}$, instead of a single threshold $r_{0}$. In our application in Section 5, we use the permutation-based test with estimator $\widehat{\lambda}^{2}_{n}$.
Table 1: Power of the three proposed statistical tests (i.e., percentage of null hypotheses rejected) at significance levels $\alpha^{*}=0.01$ and $\alpha^{*}=0.05$, across different copulas and sample sizes.
$n$ copula $\theta$ or $\rho$ $\alpha^{*}=0.01$ $\alpha^{*}=0.05$
Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{1}$ Perm. $\hat{\lambda}_{n}^{2}$ Asym. $\hat{\lambda}_{n}^{1}$
1000 Gumbel $\theta=1.1$ 5 10 7 18 26 19
$\theta=1.5$ 51 83 51 74 95 77
$\theta=2.0$ 66 96 67 86 99 87
$\theta=3.0$ 75 98 74 92 100 91
Student’s $t$ $\rho=0.1$ 2 3 2 9 10 10
$\rho=0.4$ 22 40 22 44 66 45
$\rho=0.6$ 46 77 49 71 93 74
$\rho=0.8$ 67 94 67 86 98 88
10000 Gumbel $\theta=1.1$ 18 31 18 38 57 37
$\theta=1.5$ 62 93 64 84 99 84
$\theta=2.0$ 74 98 73 90 99 90
$\theta=3.0$ 76 98 78 91 100 92
Student’s $t$ $\rho=0.1$ 2 4 3 9 13 10
$\rho=0.4$ 24 51 28 49 74 51
$\rho=0.6$ 52 87 53 76 96 75
$\rho=0.8$ 70 96 68 88 100 85